# Conditional normality and finite-state dimensions revisited

### v1

@@ 3.2 From automatic complexity to block frequencies
Assume that $A_{1}A_{2}$ with condition $B_{1}B_{2}$ has some description $P$ with respect to $M$. This means that we can read $A_{1}A_{2}$ having condition $B_{1}B_{2}$ along some path in $M$. Choose a point in this path when bits of $A_{1}$ are produced and bits of $B_{1}$ are read. (Recall that input and output bits are synchronized.) This point divides $P$ into $P_{1}P_{2}$, where $P_{1}$ is a description of $A_{1}$ with condition $B_{1}$, and $P_{2}$ is a description of $A_{2}$ with condition $B_{2}$, so we get the required bound for the sum of complexities. ∎
Remark.
This proof shows why we need to allow arbitrary initial (and final) state in the definition of automatic complexity: if we required all paths to start from some fixed initial state, then $P_{2}$ would not be a description.
@@
[…]
@@
where $\mathcal{A}$ and $\mathcal{B}$ are random variables $A_{i}$ and $B_{i}$ for uniformly distributed $i\in\{1,\ldots,u\}$, i.e., the variables considered in the definition of block entropy where $H(\mathcal{A}\mskip 0.8mu|\mskip 0.8mu\mathcal{B})$ was denoted by $H_{k,u}(\alpha\mskip 0.8mu|\mskip 0.8mu\beta)$. Combining all these inequalities and dividing by the number of bits $t$ (we assume for now that $t$ is a multiple of $k$) we get
[…]
@@
[…]
@@
[…]
Remarks.
1. Note that we have $\limsup_{k}$ in the right hand side while we had $\inf_{k}$ in the reverse inequality of the previous section. This shows that the limit over $k$ always exists and concides with infinum over $k$ (so we may use $\lim_{k}$, $\limsup_{k}$, $\liminf_{k}$, or $\inf_{k}$ in the definition of dimension in terms of block entropy).
2. One could wish to avoid using notions from Kolmogorov complexity theory in this argument. For that we could note that if every string can be a description for at most $O(1)$ strings, we can add $O(1)$ bits to encoding to get a unique code (but not a prefix-free one). To get a prefix-free code from the unique code we prepend each encoding by the prefix-free encoding of its length. This gives prefix-free code with only logarithmic overhead. (Note that we may assume without loss of generality that the description length is bounded by $k+O(1)$, so the overhead is $O(\log k)$ for $k$-bit blocks.)
@@ 3.4 Superadditivity criterion
Another byproduct of the argument above is the characterization of finite-state dimension in terms of superadditive complexity measures. Indeed, the automatic complexity function $\operatorname{\mathrm{C}}_{M}$ for each automatic description mode $M$ is superadditive and calibrated. On the other hand, for every superadditive calibrated function we can apply the same argument that worked for the automatic complexity. The only technical problem is that we cannot claim that superadditive calibrated function is an upper bound for Kolmogorov (conditional) complexity. Still it is easy to see that it is an upper bound (with $O(1)$-precision) for Kolmogorov complexity with some oracle (and the rest of the argument remains unchanged). For example, we may use the function $K$ itself as an oracle, then for every $m$ and for every condition we can enumerate all strings that have the $K$-value at most $m$ for that condition; we have $O(2^{m})$ of them and they can be encoded by programs of length $m+O(1)$ with oracle $K$. (Alternatively, we may use the combinatorial argument sketched in the remark above to avoid Kolmogorov complexity completely.)
The superadditivity criterion will be useful also for the analysis of two remaining definitions: finite-state a priori probability and gambling.
@@ 3.5 Finite-state a priori complexity
Let us show that a priori complexity definition of dimensions gives the same values as the definitions we already studied. One inequality is guaranteed by the following lemma. Recall that finite-state a priori complexity was defined in terms of labeled graphs of a special type.
@@
To prove this lemma, we need an upper bound for the value $m_{M,s}(A_{1}A_{2}\mskip 0.8mu|\mskip 0.8muB_{1}B_{2})$ for arbitrary state strings $A_{1}$, $B_{1}$ (of equal length), for arbitrary strings $A_{2},B_{2}$ (that also have equal length), and for arbitrary initial state $s$. To generate $A_{1}A_{2}$ with condition $B_{1}B_{2}$, starting from some state $s$, the probabilistic process should first generate $A_{1}$ with condition $B_{1}$ (coming to some random state $s^{\prime}$), and then generate $A_{2}$ with condition $B_{2}$ starting from $s^{\prime}$. Therefore,
@@
To get an inequality in the other direction, let us consider first the unconditional case. Consider a finite bit sequence split into several $k$-bit blocks; assume we have $N$ blocks $A_{1},\ldots,A_{N}$, each of length $k$. We want to have an upper bound for a priori complexity $\operatorname{\mathrm{KA}}_{M}(A_{1}\ldots A_{N})$ of this sequence for some finite-state probabilistic process $M$. We construct $M$ starting from some probability distribution $P$ on blocks. The probabilistic process $M$ starts from the root of a full binary tree of height $k$ and then traverses the tree generating a $k$-bit block with the probability taken from the given distribution. (For example, the root has two outgoing edges, generating bits 00 and $1$, and the probability to generate $0/1$ is the total probability of all blocks that start with $0/1$.) After $k$ bits (forming a block) are generated, the process returns to root, and the next block is generated in the same way independently from the previous ones.
@@
Remarks.
1. One could say, looking at the Shannon–Fano theorem and its proof, that probability distributions are “codes after the relaxation”: we do not require the code length $l_{i}=\log(1/p_{i})$ to be integer; so our task for $\operatorname{\mathrm{KA}}_{M}$ is easier that the corresponding task for $\operatorname{\mathrm{C}}_{M}$.
2. Due to this relaxation, we can replace the constant $1$ in the statement of the lemma by arbitrarily small positive number $\varepsilon$. Then we need to use the compactness argument: for each $Q$ we may find some $P$ (not on the boundary of the simplex of distributions) that $\varepsilon$-serves this $Q$; it has to $\varepsilon$-serve $Q$ together with some small neighborhood, and then we can find a finite cover using compactness. So the increase in block size (that was necessary to avoid overhead for codes) may be avoided now. But we already know that it is possible, so there is no reason to avoid it.
Now, using this lemma, we finish the proof in the same way as in Section 3.1. Having a finite family of graphs corresponding to distributions from the Lemma, we take a disjoint union of them. Our definition (where we take maximum probability over all initial states) guarantees that all the bounds provided by the individual processes remain valid for the disjoint union.
This was the argument for unconditional case, but we have to deal with the conditions. Our random process uses the probability distribution on blocks that depends on the condition block. So our random process uses the look-ahead bits (we need $k$ of them, where $k$ is the block size) and a finite memory that stores them. More precisely, we consider all mappings of $\mathbb{B}^{k}$ to $\mathcal{P}$ where $\mathbb{B}^{k}$ is the set of all $k$-bit strings, and $\mathcal{P}$ is the family of distributions from the Lemma. For every mapping $F$ of this type we construct a random process with look-ahead $k$ that for every condition $B$ generates bits according to $F(B)$, and then take the disjoint union of those processes for all $F$. (Technically there is a problem with the generation of the first few blocks, since the first $c$ condition bits are inaccessible, but this happens only for $O(1)$ blocks, and we can just use some fixed distribution where all blocks have positive probabilities instead of a distribution from $\mathcal{P}$ for those exceptional $O(1)$ blocks.)
Remark.
One can use a bigger class of random processes in the definition of a priori finite-state complexity. Namely, one may allow several outgoing edges with the same output letter and the same input letter but going to different vertices. (In this way the current state is not determined by input and output bits.) This wider class still gives a superadditive function $\operatorname{\mathrm{KA}}_{M}$, so the first part of the proof remains valid (and the second part becomes easier if we extend the class).
@@ 3.6 Gambling definition
The same correspondence can be explained as follows. Assume we have some measure $P$ on the Cantor space. The corresponding gambling strategy divides the initial capital in proportion $P(0){:}P(1)$ between bets on 00 and $1$. The same is done at the following stage; we use the conditional probabilities of 00 and $1$ after already known prefix to determine the next bets. Finite-state martingales correspond to output measures of finite-state random processes of the type we considered, with fixed initial state. This correspondence works as well for the conditional case.
@@
Note that a sum of martingales is a martingale, so we can consider the set of gamblers as one financial institution that makes bets. The problem is that this new combined martingale is no more a finite-state one, since the resulting proportion for the entire institution depends not only on the proportions chosen by individual gamblers, but also on their current capitals, so infinitely many ratios are possible. So to make the argument above valid for gambling, one could extend the notion of finite-state gambling by considering several gamblers with separated accounts. This notion is known under the name “$k$-account $s$-gale”.
@@
Acknowledgements. I am grateful to all the colleagues in LIRMM (Montpellier) and other places with whom I discussed the notions of normality and finite-state dimensions, especially to Ruslan Ishkuvatov and Alexander Kozachinskiy.

### v2

@@ 3.2 From automatic complexity to block frequencies
Assume that $A_{1}A_{2}$ with condition $B_{1}B_{2}$ has some description $P$ with respect to $M$. This means that we can read $A_{1}A_{2}$ having condition $B_{1}B_{2}$ along some path in $M$. Choose a point in this path when all bits of $A_{1}$ are produced and all bits of $B_{1}$ are read. (Recall that input and output bits are synchronized.) This point divides $P$ into $P_{1}P_{2}$, where $P_{1}$ is a description of $A_{1}$ with condition $B_{1}$, and $P_{2}$ is a description of $A_{2}$ with condition $B_{2}$, so we get the required bound for the sum of complexities. ∎
Remark 6.
This proof shows why we need to consider arbitrary initial (and final) state in the definition of automatic complexity: if we required all paths to start from some fixed initial state, then $P_{2}$ would not be a description.
@@
$\operatorname{\mathrm{C}}_{M}(a_{1}\ldots a_{t}\mskip 0.8mu|\mskip 0.8mub_{1}\ldots b_{t})\geqslant\sum_{i=1}^{u}\operatorname{\mathrm{K}}(A_{i}\mskip 0.8mu|\mskip 0.8muB_{i})-uO(\log k).$
@@
where $\mathcal{A}$ and $\mathcal{B}$ are random variables $A_{i}$ and $B_{i}$ for uniformly distributed $i\in\{1,\ldots,u\}$, i.e., the variables considered in the definition of block entropy, where $H(\mathcal{A}\mskip 0.8mu|\mskip 0.8mu\mathcal{B})$ was denoted by $H_{k,u}(\alpha\mskip 0.8mu|\mskip 0.8mu\beta)$. Combining all these inequalities and dividing by the number of bits $t$ (we assume for now that $t$ is a multiple of $k$) we get
$\frac{\operatorname{\mathrm{C}}_{M}(a_{1}\ldots a_{t}\mskip 0.8mu|\mskip 0.8mub_{1}\ldots b_{t})}{t}\geqslant\frac{H_{k,t/k}(\alpha\mskip 0.8mu|\mskip 0.8mu\beta)}{k}-O\left(\frac{\log k}{k}\right)$
As $t$ goes to infinity (and $k$ is fixed), the incomplete block influence becomes negligible, and for $\liminf$ and $\limsup$ it does not matter, so we forget about our assumption ($t$ is a multiple of $k$) and have
@@
for every block size $k$, and similar inequality with $\limsup$ in both sides. Then we take $\limsup_{k}$ in the right hand side, and the term $O(\log k/k)$ disappears. (Note that the left hand side does not depend on $k$.) After that the right hand side does not depend on $M$ (recall that the error term for fixed $k$ did depend on $M$), and we conclude that
@@
and this is the required inequality for two definitions of finite-state conditional dimension. The same argument works for $\limsup$.
Remark 7.
Note that we have $\limsup_{k}$ in the right hand side while we had $\inf_{k}$ in the reverse inequality of Section 3.1. This shows that the limit over $k$ always exists and coincides with infimum over $k$ (so we may use $\lim_{k}$, $\limsup_{k}$, $\liminf_{k}$, or $\inf_{k}$ in the definition of dimension in terms of block entropy).
Remark 8.
One could wish to avoid using notions from Kolmogorov complexity theory in this argument. For that we could note that if every string can be a description for at most $O(1)$ strings, we can add $O(1)$ bits to encoding to get a unique code (but not a prefix-free one). To get a prefix-free code from the unique code we prepend each encoding by the prefix-free encoding of its length. This gives prefix-free code with only logarithmic overhead. (Note that we may assume without loss of generality that the description length is bounded by $k+O(1)$, so the overhead is $O(\log k)$ for $k$-bit blocks.)
@@ 3.4 Superadditivity criterion
Another byproduct of the argument above is the characterization of finite-state dimensions in terms of superadditive complexity measures. Indeed, the automatic complexity function $\operatorname{\mathrm{C}}_{M}$ for each automatic description mode $M$ is superadditive and calibrated. On the other hand, for every superadditive calibrated function we can apply the same argument that worked for the automatic complexity. The only technical problem is that we cannot claim that superadditive calibrated function is an upper bound for Kolmogorov (conditional) complexity. Still it is easy to see that it is an upper bound (with $O(1)$-precision) for Kolmogorov complexity with some oracle (and the rest of the argument remains unchanged). For example, we may use the function $K$ (an arbitrary superadditive complexity measure given to us; nothing is assumed about its computability) as an oracle, then for every $m$ and for every condition we can enumerate all strings that have the $K$-value at most $m$ for that condition; we have $O(2^{m})$ of them and they can be encoded by programs of length $m+O(1)$ with oracle $K$. (Alternatively, we may use the combinatorial argument sketched in the Remark 8 above to avoid Kolmogorov complexity completely.)
The superadditivity criterion will be useful also for the analysis of two remaining definitions that use finite-state a priori complexity and gambling.
@@ 3.5 Finite-state a priori complexity
Let us show that a priori complexity definition of dimensions gives the same dimensions as the definitions we already studied. One inequality is guaranteed by the following lemma. Recall that finite-state a priori complexity was defined in terms of labeled graphs of a special type.
@@
To prove this lemma, we need to provide an upper bound for the value $m_{M,s}(A_{1}A_{2}\mskip 0.8mu|\mskip 0.8muB_{1}B_{2})$ for arbitrary state strings $A_{1}$, $B_{1}$ (of equal length), for arbitrary strings $A_{2},B_{2}$ (that also have equal length), and for arbitrary initial state $s$. To generate $A_{1}A_{2}$ with condition $B_{1}B_{2}$, starting from some state $s$, the probabilistic process should first generate $A_{1}$ with condition $B_{1}$ (coming to some random state $s^{\prime}$), and then generate $A_{2}$ with condition $B_{2}$ starting from $s^{\prime}$. Therefore,
@@
To get an inequality in the other direction, let us consider first the unconditional case. Consider a finite bit sequence split into several $k$-bit blocks; assume we have $N$ blocks $A_{1},\ldots,A_{N}$, each of length $k$. We want to have an upper bound for a priori complexity $\operatorname{\mathrm{KA}}_{M}(A_{1}\ldots A_{N})$ of this sequence for some finite-state probabilistic process $M$. We construct $M$ starting from some probability distribution $P$ on blocks; it will generate blocks independently according to $P$. The probabilistic process $M$ starts from the root of a full binary tree of height $k$ and then traverses the tree generating a $k$-bit block with the probability taken from $P$. (For example, the root has two outgoing edges, generating bits 0 and $1$, and the probability to generate $0/1$ is the total $P$-probability of all blocks that start with $0/1$.) After $k$ bits (forming a block) are generated, the process returns to root, and the next block is generated in the same way independently from the previous ones.
@@
Remark 9.
Recalling the Shannon–Fano theorem and its proof, one can think about probability distributions as “relaxed codes”: when the code length $l_{i}=\log(1/p_{i})$ is not required to be integer; so our task for $\operatorname{\mathrm{KA}}_{M}$ is easier that the corresponding task for $\operatorname{\mathrm{C}}_{M}$.
Remark 10.
Due to this relaxation, we can replace the constant $1$ in the statement of the lemma by arbitrarily small positive number $\varepsilon$. This strong version of the lemma can be proven using the compactness argument: for each $Q$ we may find some $P$ (not on the boundary of the simplex of distributions) that $\varepsilon$-serves this $Q$; it has to $\varepsilon$-serve $Q$ together with some small neighborhood, and then we can find a finite cover using compactness. So the increase in block size (that was necessary to avoid overhead for codes) may be avoided now. But we already know that it is possible, so there is no reason to avoid it.
Now, using this lemma, we finish the proof in the same way as in Section 3.1. Having a finite family of graphs corresponding to the distributions from the Lemma, we take a disjoint union of them. Our definition (where we take maximum probability over all initial states) guarantees that all the bounds provided by the individual processes remain valid for the disjoint union.
This was the argument for unconditional case, but we have to deal with the conditions. Our random process uses the probability distribution on blocks that depends on the condition block. So our random process uses the look-ahead bits (we need $k$ of them, where $k$ is the block size) and a finite memory that stores them. More precisely, we consider all mappings of $\mathbb{B}^{k}$ to $\mathcal{P}$ where $\mathbb{B}^{k}$ is the set of all $k$-bit strings, and $\mathcal{P}$ is the family of distributions from the Lemma. For every mapping $F$ of this type and for all possible values of the first condition block we construct a random process with look-ahead $k$ that for every condition $B$ generates bits according to $F(B)$. (Note that the first condition block is known, the second condition block is read in parallel with generating the first $k$ bits, and so on.) The we take the disjoint union of those processes for all $F$ and for all values of the first condition block.
Remark 11.
One can use a bigger class of random processes in the definition of a priori finite-state complexity. Namely, one may allow several outgoing edges with the same output letter and the same input letter but going to different vertices. (In this way the current state is not determined by input and output bits.) This wider class still gives a superadditive function $\operatorname{\mathrm{KA}}_{M}$, so the first part of the proof remains valid (and the second part becomes easier when we extend the class).
@@ 3.6 Gambling definition
The same correspondence can be explained as follows. Assume we have some measure $P$ on the Cantor space. The corresponding gambling strategy divides the initial capital in proportion P(0):P(1)P(0)\,{:}\,P(1) between bets on 0 and $1$. The same is done at the following stage; we use the conditional probabilities of 0 and $1$ after already known prefix to determine the next bets. Finite-state martingales correspond to output measures of finite-state random processes of the type we considered, with fixed initial state. This correspondence works as well for the conditional case.
@@
Note that a sum of martingales is a martingale, so we can consider the set of gamblers as one financial institution that makes bets. The problem is that this new combined martingale is no more a finite-state one, since the resulting proportion for the entire institution depends not only on the proportions chosen by individual gamblers, but also on their current capitals, so infinitely many ratios are possible. So to make the argument above valid for gambling, one could extend the notion of finite-state gambling by considering several gamblers with separated accounts. As we have noted in Section 1.2, this notion is known under the name “$k$-account $s$-gale”.
@@
  • We considered normality with respect to the uniform Bernoulli measure. Is is possible to extend this notion and its equivalent characterizations to some wider class of measures (Markov chains, or non-uniform Bernoulli measures, or some classes of product measures)?
Acknowledgements. I am grateful to all the colleagues in LIRMM (Montpellier) and other places with whom I discussed the notions of normality and finite-state dimensions, especially to Ruslan Ishkuvatov and Alexander Kozachinskiy. I am grateful to the anonymous referee (Theoretical Computer Science) for important remarks.